# RoseLoRA: Row and Column-wise Sparse Low-rank Adaptation of Pre-trained Language Model for Knowledge Editing and Fine-tuning

### v2

@@ 1 Introduction
To address this challenge, parameter-efficient fine-tuning (PEFT) methods Ding et al. (2023a); Han et al. (2024) are proposed. These method introduce a small number of learnable parameters and update only the lightweight introduced parameters during fine-tuning. Among existing methods, LoRA family Hu et al. (2021); Zhang et al. (2023); Ding et al. (2023a); Liu et al. (2024) has gained remarkable popularity because of its high efficiency and good performance. Conceptually, these LoRA methods add new low-rank matrices to model weights for fine-tuning. Unlike other PEFT methods such as Adapter (Houlsby et al., 2019), LoRA family does not modify the model architecture and is easier to incorporate.
@@ 2.1 Parameter Efficient Fine-Tuning (PEFT)
PEFT injects a small fraction of trainable parameters into pre-trained large language models (LLMs) to adapt them to downstream tasks. Prefix Tuning (Li and Liang, 2021) prepends soft tokens to the input and learns their continuous embeddings while keeping the original parameters frozen. Adapter (Houlsby et al., 2019; He et al., 2021), on the other hand, inserts lightweight bottleneck neural network modules into the transformer blocks. The third paradigm, LoRA and its variants (Hu et al., 2021; Zhang et al., 2023; Ding et al., 2023b; Dettmers et al., 2024; Li et al., 2023a; Liu et al., 2024), learns low-rank matrices to approximate the desired updates of the original model weights and has achieved state-of-the-art performance. Recently, ReFT (Wu et al., 2024) learns low-rank updates on model representations instead of weights and achieves performance comparable to LoRA with significantly fewer parameters. However, the underlying linear representation hypothesis may not hold valid (Engels et al., 2024), which greatly undermines its generalization ability. In this work, we propose an effective method to learn sparse and low-rank updates on model weights, demonstrating superior performance using as few parameters as ReFT. Recent works such as AdaLoRA (Zhang et al., 2023) and SoRA (Ding et al., 2023b) have applied pruning to LoRA to increase its computational efficiency. However, it is worth mentioning that the proposed RoseLoRA is significantly different from these methods. In particular, these works prunes to control the rank of learned model updates, but the updates are still dense in the sense that all parameters are affected, and cannot offer precise updates as RoseLoRA thereof.
@@ 3.2 Sensitivity-based Importance Score for Pruning
Importance-aware pruning Sanh et al. (2020); Han et al. (2015); Molchanov et al. (2019); Zhang et al. (2022); Li et al. (2023b) aims to identify and set redundant model weights to zero based on estimated importance scores. Parameters with high importance scores are retained, while others are set to zero. Sensitivity Sanh et al. (2020); Molchanov et al. (2019); Li et al. (2023b) is a popular importance metric that measures the approximate change in training loss when setting a parameter to zero. Formally, the sensitivity with respect to weight $\bm{W}_{ij}$ is defined by the product of the weight and its corresponding gradient:
@@ 4.2 Optimization
Here, $\tau^{(t)}$ is the budget of the percentage of remaining parameters at the $t$-iteration. To enable the optimization to be more stable, we decrease the number of $\tau^{(t)}$ gradually following the cubic strategy Li et al. (2023b):
$\tau^{(t)}=\left\{\begin{aligned} &1,&1\leq t\leq t_{i},\\ &\tau+(1-\tau)(1-\frac{t-t_{i}}{t_{f}-t_{i}})^{3},&t_{i}\leq t\leq t_{f},\\ &\tau,&t_{f}\leq t\leq T,\end{aligned}\right.$
@@ 5.2 Performance Comparison
Table 1: Performance comparison of LLaMA 2-7b-chat against existing knowledge editing methods on four knowledge editing datasets. Results marked with $\heartsuit$” are taken from Zhang et al. (2024). ”AVG” means the average of edit success, locality, portability, and fluency. Because fluency is not at the same magnitude as other metrics, we leverage ”fluency/10” when computing AVG values.
@@
Table 2: Accuracy comparison of LLaMA-7B against PEFT baselines on eight commonsense reasoning datasets. Results marked with $\heartsuit$” are taken from Liu et al. (2024). ”AVG” means the average accuracy of all datasets. For RoseLoRA, Params (%) is calculated by dividing the number of final low-rank matrices parameters by the number of parameters of the base LMs (number of low-rank matrix parameters times sparsity).
@@
Table 3: Accuracy comparison of LLaMA-7B against PEFT baselines on four arithmetic reasoning datasets. Results marked with $\heartsuit$” are taken from Hu et al. (2023). ”AVG” means the average accuracy of all datasets.
@@
Table 4: Performance comparison of LLaMA-2 7B on instruction tuning task on Alpaca-Eval v1.0. We compute the win-rate against text-davinci-003 using GPT-4 as the annotator. Results marked with $\heartsuit$” are taken from Wu et al. (2024).
@@
Table 5: Accuracy comparison of RoBERTa-large against PEFT baselines on the GLUE benchmark. Results marked with $\heartsuit$” are taken from Wu et al. (2023). ”AVG” means the average accuracy of all datasets.
@@ 5.3 Forgetting Test
Table 6: Accuracy of fine-tuned models on TriviaQA (knowledge reasoning), MMLU (general knowledge), and ARC-c (commonsense reasoning) dataset. ”AVG” is the average accuracy of Humanities, Social Sciences, STEM, and Other fields on MMLU. The evaluation is conducted with Lm-Evaluation-Harness (Gao et al., 2023).
@@ Appendix A Proof of Proposition 1
WikiData recent 2e-4 4 1 0.95 0.8 3e-3 ”up_proj”, ”down_proj”, ”gate_proj”
WikiData counterfact 2e-4 4 1 0.95 3e-3 ”up_proj”, ”down_proj”, ”gate_proj”
ZsRE 2e-4 4 1 0.95 3e-3 ”up_proj”, ”down_proj”, ”gate_proj”
WikiBio 2e-4 4 1 0.95 3e-3 ”up_proj”, ”down_proj”, ”gate_proj”
Commonsense170K 2e-4 32 8 0.865 - ”q_proj”,”v_proj”
Math10K 3e-4 32 32 0.865 - ”q_proj”,”v_proj”
Instruction tuning 3e-4 32 32 0.85 - ”q_proj”,”v_proj”
@@
CoLA Matthews corr 2e-4 6 16 0.95 0.8 ”query”, ”key”, ”value”, ”output.dense”, ”intermediate.dense”
@@ Appendix B Datasets, Metrics and Hyper-parameters
  • Commonsense reasoning contains of eight datasets, including BoolQ [Clark et al., 2019], PIQA [Bisk et al., 2020], SIQA [Sap et al., 2019], HellaSwag [Zellers et al., 2019], WinoGrande [Sakaguchi et al., 2021], ARC-e, ARC-c [Clark et al., 2018], and OBQA [Mihaylov et al., 2018]. These tasks are multiple choice problems. Following Hu et al. [2023], Wu et al. [2024], we fine-tune the LLM on a combined training dataset named Commonsense170K of these tasks and evaluate the Accuracy on individual test sets.
  • Arithmetic reasoning consists of four math reasoning datasets: AQuA [Ling et al., 2017], GSM8K [Cobbe et al., 2021], MAWPS [Koncel-Kedziorski et al., 2016], and SVAMP [Patel et al., 2021]. Models need to generate correct answers and we use Accuracy as the evaluation metric following Hu et al. [2023] as well. Again, we replicate the setup in Wu et al. [2024] and fine-tune the models on the combined training data named Math10K of the four tasks.
  • Instruction-following measures if the model can follow human instructions. Same as before, we follow Hu et al. [2023], Wu et al. [2024] and use Ultrafeedback [Cui et al., 2023] as the training data, and evaluate the model performance by Alpaca-Eval v1.0 [Li et al., 2023c].
  • Natural language understanding consists of eight datasets from the GLUE benchmark [Wang et al., 2018]. We adopt the evaluation metrics and setups from Wu et al. [2023].
@@ Appendix C Implementation of Knowledge Editing
To enable the minimal modification of the LLM, following Zhang et al. [2024], we add one $\ell_{2}$ norm on the low-rank matrices:

### v3

@@ Abstract
RoseLoRA: Row and Column-wise Sparse Low-rank Adaptation of Pre-trained Language Model for Knowledge Editing and Fine-tuning
Haoyu Wang, Tianci Liu§, Ruirui Li, Monica Xiao Cheng, Tuo Zhao, and Jing Gao§ SUNY Albany, Albany, NY, USA §Purdue University, West Lafayette, IN, USA Georgia Institute of Technology, Atlanta, GA, USA Amazon, Palo Alto, CA, USA hwang28@albany.edu, §{liu3351,jinggao}@purdue.edu, tourzhao@gatech.edu , {ruirul,chengxca}@amazon.com
@@ 1 Introduction
To address this challenge, parameter-efficient fine-tuning (PEFT) methods Ding et al. (2023b); Han et al. (2024) are proposed. These method introduce a small number of learnable parameters and update only the lightweight introduced parameters during fine-tuning. Among existing methods, LoRA family Hu et al. (2021); Zhang et al. (2023); Ding et al. (2023b); Liu et al. (2024) has gained remarkable popularity because of its high efficiency and good performance. Conceptually, these LoRA methods add new low-rank matrices to model weights for fine-tuning. Unlike other PEFT methods such as Adapter (Houlsby et al., 2019), LoRA family does not modify the model architecture and is easier to incorporate.
@@ 2.1 Parameter Efficient Fine-Tuning (PEFT)
PEFT injects a small fraction of trainable parameters into pre-trained large language models (LLMs) to adapt them to downstream tasks. Prefix Tuning (Li and Liang, 2021) prepends soft tokens to the input and learns their continuous embeddings while keeping the original parameters frozen. Adapter (Houlsby et al., 2019; He et al., 2021), on the other hand, inserts lightweight bottleneck neural network modules into the transformer blocks. The third paradigm, LoRA and its variants (Hu et al., 2021; Zhang et al., 2023; Ding et al., 2023a; Dettmers et al., 2024; Li et al., 2023b; Liu et al., 2024), learns low-rank matrices to approximate the desired updates of the original model weights and has achieved state-of-the-art performance. Recently, ReFT (Wu et al., 2024) learns low-rank updates on model representations instead of weights and achieves performance comparable to LoRA with significantly fewer parameters. However, the underlying linear representation hypothesis may not hold valid (Engels et al., 2024), which greatly undermines its generalization ability. In this work, we propose an effective method to learn sparse and low-rank updates on model weights, demonstrating superior performance using as few parameters as ReFT. Recent works such as AdaLoRA (Zhang et al., 2023) and SoRA (Ding et al., 2023a) have applied pruning to LoRA to increase its computational efficiency. However, it is worth mentioning that the proposed RoseLoRA is significantly different from these methods. In particular, these works prunes to control the rank of learned model updates, but the updates are still dense in the sense that all parameters are affected, and cannot offer precise updates as RoseLoRA thereof.
@@ 3.2 Sensitivity-based Importance Score for Pruning
Importance-aware pruning Sanh et al. (2020); Han et al. (2015); Molchanov et al. (2019); Zhang et al. (2022); Li et al. (2023c) aims to identify and set redundant model weights to zero based on estimated importance scores. Parameters with high importance scores are retained, while others are set to zero. Sensitivity Sanh et al. (2020); Molchanov et al. (2019); Li et al. (2023c) is a popular importance metric that measures the approximate change in training loss when setting a parameter to zero. Formally, the sensitivity with respect to weight $\bm{W}_{ij}$ is defined by the product of the weight and its corresponding gradient:
@@ 4.2 Optimization
Here, $\tau^{(t)}$ is the budget of the percentage of remaining parameters at the $t$-iteration. To enable the optimization to be more stable, we decrease the number of $\tau^{(t)}$ gradually following the cubic strategy Li et al. (2023c):
$\tau^{(t)}=\left\{\begin{aligned} &1,&1\leq t\leq t_{i},\\ &\tau+(1-\tau)\left(1-\frac{t-t_{i}}{t_{f}-t_{i}}\right)^{3},&t_{i}\leq t\leq t_{f},\\ &\tau,&t_{f}\leq t\leq T,\end{aligned}\right.$
@@ 5.2 Performance Comparison
Table 1: Performance comparison of LLaMA-7b-chat against existing knowledge editing methods on four knowledge editing datasets. Results marked with $\heartsuit$" are taken from Zhang et al. (2024). "AVG" means the average of edit success, locality, portability, and fluency. Because fluency is not at the same magnitude as other metrics, we leverage "fluency/10" when computing AVG values.
@@
Table 2: Accuracy comparison of LLaMA-7B against PEFT baselines on eight commonsense reasoning datasets. Results marked with $\heartsuit$" are taken from Liu et al. (2024). "AVG" means the average accuracy of all datasets. For RoseLoRA, Params (%) is calculated by dividing the number of final low-rank matrices parameters by the number of parameters of the base LMs (number of low-rank matrix parameters times sparsity).
@@
Table 3: Accuracy comparison of LLaMA-7B against PEFT baselines on four arithmetic reasoning datasets. Results marked with $\heartsuit$" are taken from Hu et al. (2023). "AVG" means the average accuracy of all datasets.
@@
Table 4: Performance comparison of LLaMA-2 7B on instruction tuning task on Alpaca-Eval v1.0. We compute the win-rate against text-davinci-003 using GPT-4 as the annotator. Results marked with $\heartsuit$" are taken from Wu et al. (2024).
@@
Table 5: Accuracy comparison of RoBERTa-large against PEFT baselines on the GLUE benchmark. Results marked with $\heartsuit$" are taken from Wu et al. (2023). "AVG" means the average accuracy of all datasets.
@@ 5.3 Forgetting Test
Table 6: Accuracy of fine-tuned models on TriviaQA (knowledge reasoning), MMLU (general knowledge), and ARC-c (commonsense reasoning) dataset. "AVG" is the average accuracy of Humanities, Social Sciences, STEM, and Other fields on MMLU. The evaluation is conducted with Lm-Evaluation-Harness (Gao et al., 2023).
@@ 6 Conclusion
Limitations
The proposed RoseLoRA framework introduces a hyper-parameter $\beta$ to smooth the sensitivity estimation, which might require additional effort to tune. Fortunately, we observe that the model performance is not sensitive to the hyper-parameter and we set it to a fixed value to achieve good performance in this paper.
Acknowledgement
This work is supported in part by the US National Science Foundation under grant NSF IIS-1747614 and NSF IIS-2141037. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Proof of Proposition 1
WikiData recent 2e-4 4 1 0.95 0.8 3e-3 "up_proj", "down_proj", "gate_proj"
WikiData counterfact 2e-4 4 1 0.95 3e-3 "up_proj", "down_proj", "gate_proj"
ZsRE 2e-4 4 1 0.95 3e-3 "up_proj", "down_proj", "gate_proj"
WikiBio 2e-4 4 1 0.95 3e-3 "up_proj", "down_proj", "gate_proj"
Commonsense170K 2e-4 32 8 0.865 - "q_proj","v_proj"
Math10K 3e-4 32 32 0.865 - "q_proj","v_proj"
Instruction tuning 3e-4 32 32 0.85 - "q_proj","v_proj"
@@
CoLA Matthews corr 2e-4 6 16 0.95 0.8 "query", "key", "value", "output.dense", "intermediate.dense"
@@ Appendix B Datasets, Metrics and Hyper-parameters
  • Commonsense reasoning contains of eight datasets, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). These tasks are multiple choice problems. Following Hu et al. (2023); Wu et al. (2024), we fine-tune the LLM on a combined training dataset named Commonsense170K of these tasks and evaluate the Accuracy on individual test sets.
  • Arithmetic reasoning consists of four math reasoning datasets: AQuA (Ling et al., 2017), GSM8K (Cobbe et al., 2021), MAWPS (Koncel-Kedziorski et al., 2016), and SVAMP (Patel et al., 2021). Models need to generate correct answers and we use Accuracy as the evaluation metric following Hu et al. (2023) as well. Again, we replicate the setup in Wu et al. (2024) and fine-tune the models on the combined training data named Math10K of the four tasks.
  • Instruction-following measures if the model can follow human instructions. Same as before, we follow Hu et al. (2023); Wu et al. (2024) and use Ultrafeedback (Cui et al., 2023) as the training data, and evaluate the model performance by Alpaca-Eval v1.0 (Li et al., 2023a).
  • Natural language understanding consists of eight datasets from the GLUE benchmark (Wang et al., 2018). We adopt the evaluation metrics and setups from Wu et al. (2023).
@@ Appendix C Implementation of Knowledge Editing
To enable the minimal modification of the LLM, following Zhang et al. (2024), we add one $\ell_{2}$ norm on the low-rank matrices: